OpenStudy (anonymous):

How large an average force is required to stop a 1400-kg car in 5.0 s if the car’s initial speed is 25 m/s? a 2000 N b 3500 N c 9000 N d 7 000N

OpenStudy (anonymous):

Avg Force = (mass)(change in acceleration) = (1400)(0-25)/5 = - 1400*5 = -7000 N " - "sign indicates that force acts opposite to the direction of motion of a car. Ans d)